OpenStudy (bossimbacon):

Complete the equations of the system in slope-intercept form. Use a decimal for the slope if necessary. Line 1 x y -3 6 -2 5 Line 2 x y 0 5 2 1 Line 1: y =____ x +____ Line 2: y =____ x +____

OpenStudy (bossimbacon):

@Demonx341

OpenStudy (anonymous):

line 1: -3x + 7= 4x and -2x + 5= 3x line 2: 0x + 5 = 5x and 2x + 1 = 3x

OpenStudy (bossimbacon):

@Demonx341 wait i don't understand

OpenStudy (anonymous):

just take the x out and put it at the answer. is what x means. it is a veriable that is not given so you just move it to the end. @Bossimbacon

OpenStudy (bossimbacon):

ooo thank you :)

OpenStudy (bossimbacon):

whoopsies it was a typo can you redo it:( it was -3 7 not -3 and 6 my bad. @Demonx341